# Compatibility of Kazhdan and Brauer Homomorphism

### v2

@@ 6.3. Compatibility of Brauer homomorphism with linkage
Let $\Xi$ be an irreducible $\sigma$-fixed representation of $G$, and let $T:\Xi\rightarrow\Xi\circ\sigma$ be an isomorphism such that $T^{l}={\rm id}$. For any $\sigma$-plain compact open subgroup $K$ of $G$, the inclusion map $\Xi^{K}\hookrightarrow\Xi$ is $\sigma$-invariant, and it induces a map $\widehat{H}^{i}(\Xi^{K})\rightarrow\widehat{H}^{i}(\Xi)$, taking values in the $\mathcal{H}(G^{\sigma},K^{\sigma})$-module $\widehat{H}^{i}(\Xi)^{K^{\sigma}}$. It is shown in [TV16, Section 6.2] that for all $h\in\mathcal{H}(G,K)^{\sigma}$, the Brauer homomorphism $\overline{\rm Br}$ induces the commutative diagram
@@
and this induces the compatibility of Brauer homomorphism with linkage in the sense that, for a representation $\rho$ of $G^{\sigma}$ to be linked with $\Xi$, it is equivalent to say that there exists a composition factor $N$ of the $\mathcal{H}(G,K)^{\sigma}$-module $\widehat{H}^{i}(\Xi^{K})$ such that
$\mathcal{H}(G^{\sigma},K^{\sigma})\otimes_{\mathcal{H}(G,K)^{\sigma}}N\simeq(\rho^{(l)})^{K^{\sigma}}.$ (6.3)
@@ 6.3.1. Compatibility of Kazhdan isomorphism with linkage
We first recall the notations and set up under which we study this compatibility. Let $F$ and $F^{\prime}$ be the non-Archimedean $m$-close local fields with same residue characteristic $p$. Let $E$ be the finite Galois extension of $F$ of prime degree $l$ with $l\neq p$. By Lemma 4.2, there exists a finite Galois extension $E^{\prime}/F^{\prime}$ with same degree of extension $l$. As before, we fix the generators $\sigma$ and $\sigma^{\prime}$ of the Galois groups ${\rm Gal}(E/F)$ and ${\rm Gal}(E^{\prime}/F^{\prime})$ respectively.
Let G be a split connected reductive group defined over $\mathbb{Z}$. Let $G_{E}$ (resp. $G_{F}$) be the $E$ (resp. $F$)-points of G. Let $K_{E}$ and $K_{F}$ be the $m$-th congruence subgroups of $G_{E}$ and $G_{F}$ respectively (see Section 4). Then we have
${\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}}={\rm\overline{Br}^{\prime}}\circ{\rm Kaz}_{em}^{E},$ (6.4)
where $e\in\{1,l\}$. It is shown in [Gan15, Section 2.3] that there are bijections
$\big{\{}\rho_{E}\in{\rm Irr}(G_{E}):\rho_{E}^{K_{E}}\neq 0\big{\}}\longrightarrow\big{\{}\rho_{E^{\prime}}\in{\rm Irr}(G_{E^{\prime}}):\rho_{E^{\prime}}^{K_{E^{\prime}}}\neq 0\big{\}}$ (6.5)
@@
$\big{\{}\rho_{F}\in{\rm Irr}(G_{F}):\rho_{F}^{K_{F}}\neq 0\big{\}}\longrightarrow\big{\{}\rho_{F^{\prime}}\in{\rm Irr}(G_{F^{\prime}}):\rho_{F^{\prime}}^{K_{F^{\prime}}}\neq 0\big{\}},$ (6.6)
induced by the Kazhdan isomorphisms ${\rm Kaz}_{em}^{E}$ and ${\rm Kaz}_{m}^{F}$ respectively. With these notations and terminologies, we now prove
@@
Let $(\rho_{E},V_{E})$ be an irreducible $\sigma$-fixed representation of $G_{E}$ with $V_{E}^{K_{E}}\neq 0$, and let $(\rho_{F},V_{F})$ be an irreducible representation of $G_{F}$ with $V_{F}^{K_{F}}\neq 0$. Let $(\rho_{E^{\prime}},V_{E^{\prime}})$ and $(\rho_{F^{\prime}},V_{F^{\prime}})$ be the corresponding objects under the bijections (6.5) and (6.6). Then $\rho_{F}$ is linked with $\rho_{E}$ if and only if $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$.
@@
The proof is divided into two parts. In the first part, we show that the irreducible representation $\rho_{E^{\prime}}$, obtained from $\rho_{E}$ under the bijection (6.5), is $\sigma^{\prime}$-fixed. The second part shows that the linkage between $\rho_{F}$ and $\rho_{E}$ implies the linkage between $\rho_{F^{\prime}}$ and $\rho_{E^{\prime}}$, and vice versa.
6.3.2.
First recall that the space $V_{E}^{K_{E}}$ is an irreducible $\mathcal{H}(G_{E},K_{E})$-module, and it can be regarded as $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$-module via Kazhdan isomorphism. Consider the $\mathcal{H}(G_{E^{\prime}})$ (For definition, see [Vig96, Chapter 1])-module
$N^{\prime}=\mathcal{H}(G_{E^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})}V_{E}^{K_{E}}.$
Note that the action of $\langle\sigma^{\prime}\rangle$ on $\mathcal{H}(G_{E^{\prime}})$ induces an action of $\langle\sigma^{\prime}\rangle$ on $N^{\prime}$, given by
@@
for $f^{\prime}\in\mathcal{H}(G_{E^{\prime}})$ and $v\in V_{E}^{k_{E}}$. Following the construction of [Vig96, Section 4, Chapter 1], the $\mathcal{H}(G_{E^{\prime}})$-module $N^{\prime}$ gives the irreducible representation $\rho_{E^{\prime}}$, defined as
@@
for $g^{\prime}\in G_{E^{\prime}},x^{\prime}\in V_{E^{\prime}}$. Here, *** denotes the action of the Hecke algebra $\mathcal{H}(G_{E^{\prime}})$ on the space $V_{E^{\prime}}$ and $1_{g^{\prime}K_{E^{\prime}}}$ is the characteristic function of the left coset $g^{\prime}K_{E^{\prime}}$. Since $\rho_{E}$ is $\sigma$-fixed, there exists an isomorphism $T_{\sigma}:\rho_{E}\rightarrow\rho_{E}\circ\sigma$. Then the map $T_{\sigma^{\prime}}:N^{\prime}\rightarrow N^{\prime}$, defined by
@@
induces an isomorphism $\rho_{E^{\prime}}\xrightarrow{\sim}\rho_{E^{\prime}}\circ\sigma^{\prime}$, and this shows that the representation $\rho_{E^{\prime}}$ is $\sigma^{\prime}$-fixed.
6.3.3.
Now, suppose that the representation $\rho_{F}$ is linked with $\rho_{E}$. Then the $\mathcal{H}(G_{F},K_{F})$-module $(V_{F}^{(l)})^{K_{F}}$ is a composition factor of $\widehat{H}^{i}(V_{E})^{K_{F}}$, for some $i\in\{0,1\}$. If we regard these as $\mathcal{H}(G_{E},K_{E})^{\sigma}$-modules via the Brauer homomorphism $\overline{\rm Br}$, then the formulation (6.3) gives a composition factor $M$ of $\widehat{H}^{i}(V_{E}^{K_{E}})$ such that
$\mathcal{H}(G_{F},K_{F})\otimes_{\mathcal{H}(G_{E},K_{E})^{\sigma}}M\simeq(V_{F}^{(l)})^{K_{F}}.$ (6.7)
@@
Let $M^{\prime}$ be the image of $M$ under the map $\Phi$. Then $M^{\prime}$ is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}}^{K_{E^{\prime}}})$, and the $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$-module
$\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}M^{\prime}$
is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}})^{K_{F^{\prime}}}$. We also have an isomorphism $\varphi:V_{F}^{K_{F}}\rightarrow V_{F^{\prime}}^{K_{F^{\prime}}}$, which is equivariant under the action of $\mathcal{H}(G_{F},K_{F})$ and hence $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$ via the Kazhdan isomorphism ${\rm Kaz}_{m}^{F}$. Using the isomorphisms $\Phi$, $\varphi$ and the relation (6.4), it follows from (6.7) that
$\mathcal{H}(G_{F,},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}M^{\prime}\simeq(V_{F^{\prime}}^{(l)})^{K_{F^{\prime}}}.$
This shows that the representation $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$. A similar argument also establishes the other implication, namely that the representation $\rho_{F}$ is linked with $\rho_{E}$ if $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$. This completes the proof. ∎
Acknowledgements. The author is grateful to Santosh Nadimpalli for communicating Proposition 6.2, and for his helpful comments during the work.

### v4

@@ 6.4. Compatibility of Brauer homomorphism with linkage
Let $\Xi$ be an irreducible $\sigma$-fixed representation of $G$, and let $T:\Xi\rightarrow\Xi\circ\sigma$ be an isomorphism such that $T^{l}={\rm id}$. For any $\sigma$-plain compact open subgroup $K$ of $G$, the inclusion map $\Xi^{K}\hookrightarrow\Xi$ is $\sigma$-invariant and it induces a map $\widehat{H}^{i}(\Xi^{K})\rightarrow\widehat{H}^{i}(\Xi)$, taking values in the $\mathcal{H}(G^{\sigma},K^{\sigma})$-module $\widehat{H}^{i}(\Xi)^{K^{\sigma}}$. It is shown in [TV16, Section 6.2] that, for all $h\in\mathcal{H}(G,K)^{\sigma}$, the Brauer homomorphism $\overline{\rm Br}$ induces the following commutative diagram
@@
and it induces the compatibility of Brauer homomorphism with linkage in the sense that, for a representation $\rho$ of $G^{\sigma}$ to be linked with $\Xi$, it is equivalent to say that there exists a composition factor $\mathcal{N}$ of the $\mathcal{H}(G,K)^{\sigma}$-module $\widehat{H}^{i}(\Xi^{K})$ such that
$\mathcal{H}(G^{\sigma},K^{\sigma})\otimes_{\mathcal{H}(G,K)^{\sigma}}\mathcal{N}\simeq(\rho^{(l)})^{K^{\sigma}}.$ (6.6)
@@ 6.4.1. Compatibility of Kazhdan isomorphism with linkage
Let $F$ and $F^{\prime}$ be two non-Archimedean $m$-close local fields with the same residue characteristic $p$. Let $E$ be a finite Galois extension of $F$ of prime degree $l$ with $l\neq p$. By Lemma 4.6, we have a finite Galois extension $E^{\prime}/F^{\prime}$ of degree $l$ such that $E^{\prime}$ is $em$-close to $E$, where $e=1$ if $E/F$ is unramified and $e=l$ if $E/F$ is totally ramified. Fix two generators $\sigma$ and $\sigma^{\prime}$ of the groups ${\rm Gal}(E/F)$ and ${\rm Gal}(E^{\prime}/F^{\prime})$ respectively.
Let G be a split connected reductive group defined over $\mathbb{Z}$. Let $K_{E}$ (resp. $K_{F}$) be the congruence subgroup of $G_{E}$ (resp. $G_{F}$) of level $m$ (resp. $em$) (see Section 4 for the definition). Then, from Theorem 5.1, we get the following relation:
${\rm Kaz}_{m}^{F}\circ{\rm\overline{Br}}={\rm\overline{Br}^{\prime}}\circ{\rm Kaz}_{em}^{E}.$ (6.7)
Moreover, we have the following bijections (see [Gan15, Section 2.3])
$\big{\{}(\rho_{E},V_{E})\in{\rm Irr}(G_{E}):V_{E}^{K_{E}}\neq 0\big{\}}\longrightarrow\big{\{}(\rho_{E^{\prime}},V_{E^{\prime}})\in{\rm Irr}(G_{E^{\prime}}):V_{E^{\prime}}^{K_{E^{\prime}}}\neq 0\big{\}}$ (6.8)
@@
$\big{\{}(\rho_{F},V_{F})\in{\rm Irr}(G_{F}):V_{F}^{K_{F}}\neq 0\big{\}}\longrightarrow\big{\{}(\rho_{F^{\prime}},V_{F^{\prime}})\in{\rm Irr}(G_{F^{\prime}}):V_{F^{\prime}}^{K_{F^{\prime}}}\neq 0\big{\}},$ (6.9)
induced by the isomorphisms ${\rm Kaz}_{em}^{E}$ and ${\rm Kaz}_{m}^{F}$, respectively. With these notations, we now prove
@@
Let $(\rho_{E},V_{E})$ be an irreducible $\sigma$-fixed representation of $G_{E}$ with $V_{E}^{K_{E}}\neq 0$, and let $(\rho_{F},V_{F})$ be an irreducible representation of $G_{F}$ with $V_{F}^{K_{F}}\neq 0$. Let $(\rho_{E^{\prime}},V_{E^{\prime}})$ and $(\rho_{F^{\prime}},V_{F^{\prime}})$ be the corresponding objects under the bijections (6.8) and (6.9). Then, the irreducible representation $\rho_{E^{\prime}}$ of $G_{E^{\prime}}$ is $\sigma^{\prime}$-fixed, and $\rho_{F}^{\prime}$ is linked with $\rho_{E}^{\prime}$ if and only if $\rho_{F}$ is linked with $\rho_{E}$.
@@
We divide the proof into two parts. The first part proves that the irreducible representation $\rho_{E^{\prime}}$ is $\sigma^{\prime}$-fixed. In the second part, we show that the linkage between $\rho_{F}$ and $\rho_{E}$ implies the linkage between $\rho_{F^{\prime}}$ and $\rho_{E^{\prime}}$, and vice versa.
6.4.2.
Recall that the space of $K_{E}$-fixed vectors $V_{E}^{K_{E}}$ is a simple $\mathcal{H}(G_{E},K_{E})$-module, and it can be regarded as a $\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})$-module via the Kazhdan isomorphism ${\rm Kaz}_{em}^{E}$. Consider the $\mathcal{H}(G_{E^{\prime}})$-module
$\mathcal{N}^{\prime}=\mathcal{H}(G_{E^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})}V_{E}^{K_{E}}.$
Note that the action of $\langle\sigma^{\prime}\rangle$ on $\mathcal{H}(G_{E^{\prime}})$ induces an action of $\langle\sigma^{\prime}\rangle$ on $\mathcal{N}^{\prime}$, by setting
@@
for all $f^{\prime}\in\mathcal{H}(G_{E^{\prime}})$ and $v\in V_{E}^{K_{E}}$. Following the construction of [Vig96, Chapter I, Subsection 4.4], the $\mathcal{H}(G_{E^{\prime}})$-module $\mathcal{N}^{\prime}$ gives the irreducible representation $(\rho_{E^{\prime}},V_{E^{\prime}})$, defined as
@@
for $g^{\prime}\in G_{E^{\prime}},x^{\prime}\in V_{E^{\prime}}$. Here * denotes the action of the Hecke algebra $\mathcal{H}(G_{E^{\prime}})$ on the space $V_{E^{\prime}}$ and $1_{g^{\prime}K_{E^{\prime}}}$ is the characteristic function of the left coset $g^{\prime}K_{E^{\prime}}$. Since $\rho_{E}$ is $\sigma$-fixed, there exists an isomorphism $T_{\sigma}:\rho_{E}\rightarrow\rho_{E}\circ\sigma$. Then the map $T_{\sigma^{\prime}}:\mathcal{N}^{\prime}\rightarrow\mathcal{N}^{\prime}$, defined by
@@
induces an isomorphism between $\rho_{E^{\prime}}$ and $\rho_{E^{\prime}}\circ\sigma^{\prime}$. Hence, $\rho_{E^{\prime}}$ is $\sigma^{\prime}$-fixed.
6.4.3.
Suppose the representation $\rho_{F}$ is linked with $\rho_{E}$. Then, the space of $K_{F}$-fixed vectors $(V_{F}^{(l)})^{K_{F}}$, considered as $\mathcal{H}(G_{F},K_{F})$-module, is a composition factor of $\widehat{H}^{i}(V_{E})^{K_{F}}$, for some $i$. If we regard these as $\mathcal{H}(G_{E},K_{E})^{\sigma}$-modules via the Brauer homomorphism $\overline{\rm Br}$, then the formulation (6.6) gives a composition factor $\mathcal{M}$ of $\widehat{H}^{i}(V_{E}^{K_{E}})$ such that
$\mathcal{H}(G_{F},K_{F})\otimes_{\mathcal{H}(G_{E},K_{E})^{\sigma}}\mathcal{M}\simeq(V_{F}^{(l)})^{K_{F}}.$ (6.10)
@@
Let $\mathcal{M}^{\prime}$ be the image of $\mathcal{M}$ under $\Phi$. Then $\mathcal{M}^{\prime}$ is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}}^{K_{E^{\prime}}})$, and the $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$-module
$\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}\mathcal{M}^{\prime}$
is a composition factor of $\widehat{H}^{i}(V_{E^{\prime}})^{K_{F^{\prime}}}$. We also have an isomorphism $\varphi:V_{F}^{K_{F}}\rightarrow V_{F^{\prime}}^{K_{F^{\prime}}}$, which is equivariant under the action of $\mathcal{H}(G_{F},K_{F})$ and hence of $\mathcal{H}(G_{F^{\prime}},K_{F^{\prime}})$ via Kazhdan isomorphism ${\rm Kaz}_{m}^{F}$. Using the isomorphisms $\Phi$, $\varphi$ and the relation (6.7), it follows from (6.10) that
$\mathcal{H}(G_{F,},K_{F^{\prime}})\otimes_{\mathcal{H}(G_{E^{\prime}},K_{E^{\prime}})^{\sigma^{\prime}}}\mathcal{M}^{\prime}\simeq(V_{F^{\prime}}^{(l)})^{K_{F^{\prime}}}.$
This implies that the representation $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$. A similar argument also shows that the representation $\rho_{F}$ is linked with $\rho_{E}$ if $\rho_{F^{\prime}}$ is linked with $\rho_{E^{\prime}}$. Hence, the theorem. ∎
Acknowledgements. The author expresses his deep gratitude to the anonymous referee for careful reading of the manuscript and for various suggestions. The author is grateful to Santosh Nadimpalli for communicating Proposition 6.4, and for his helpful comments during the work. The author would also like to thank Radhika Ganapathy for her valuable comments and suggestions.